# Learning nonnegative matrix factorizations from compressed data

### v1

@@ 4.1. General convergence for sketched multiplicative updates
4.2. Multiplicative updates for solving regularized compressed problems
Now, we demonstrate how the general framework (4.1) applies to the compressed problems from Section 3.
Corollary 4.3 (Two-sided updates).
Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be a nonnegative matrix and ${\mathbf{A}}_{1}\in{\mathbb{R}}^{k\times m}$, ${\mathbf{A}}_{2}\in{\mathbb{R}}^{n\times k}$ are generic conpression matrices such that ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}$ is invertible. Form ${\mathbf{Q}}_{1}\in{\mathbb{R}}^{m\times r}$ and ${\mathbf{Q}}_{2}\in{\mathbb{R}}^{n\times r}$, the matrices which columns form orthonormal bases of the column space of ${\mathbf{X}}{\mathbf{A}}_{2}$ and the row space of ${\mathbf{A}}_{1}{\mathbf{X}}$ respectively. For any $\lambda_{1},\lambda_{2}\geq 0$, if
$\sigma_{1}$ $\geq\max\{\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}\right)_{-}\},\max\{\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}+\lambda_{1}({\mathbf{I}}-{\mathbf{Q}}_{1}{\mathbf{Q}}_{1}^{T})\right)_{-}\},$
$\sigma_{2}$ $\geq\max\{\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}\right)_{-}\},\max\{\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}+\lambda_{2}({\mathbf{I}}-{\mathbf{Q}}_{2}{\mathbf{Q}}_{2}^{T})\right)_{-}\},$
where $\max$ is taken entry-wise, then the objective
@@ 4.2. Multiplicative updates for solving regularized compressed problems
$+\lambda_{1}\|P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}+\lambda_{2}\|{\mathbf{U}}{\mathbf{V}}^{T}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}\|^{2}_{F}$
@@
Note that Theorem 3.1 and Corollary 3.12 claim that the optimal solution for (4.5) is the optimal solution for the uncompressed NMF problem if ${\mathbf{X}}$ has exactly nonnegative decomposition of the rank at most $k$.
@@
Consider the setting of Theorem 4.1, with the matrices
@@
$\{{\mathbf{B}}_{j}\}_{j=1,2,3}=\{{\mathbf{A}}_{2},\sqrt{\sigma_{2}}\mathbf{1}_{n},\sqrt{\lambda_{1}}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}\}=\{{\mathbf{A}}_{2},\sqrt{\sigma_{2}}\mathbf{1}_{n},\sqrt{\lambda_{1}}({\mathbf{I}}-{\mathbf{Q}}_{1}{\mathbf{Q}}_{1}^{T})\},$
@@
Clearly, we have that the matrices ${\mathbf{X}}^{({\mathbf{A}})}_{i}$ and ${\mathbf{X}}^{({\mathbf{B}})}_{i}$ are nonnegative. Then to apply Theorem 4.1, we must check that $\sum_{i}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}$, $\sum_{i}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}{\mathbf{X}}^{({\mathbf{A}})}_{i}$, $\sum_{j}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T}$ and $\sum_{j}{\mathbf{X}}^{({\mathbf{B}})}_{j}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T}$ are entry-wise nonnegative. First, we calculate
@@
$\sigma_{1}\geq\max\{\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}+\lambda_{1}({\mathbf{I}}-{\mathbf{Q}}_{1}{\mathbf{Q}}_{1}^{T})\right)_{-}\},\;\sigma_{2}\geq\max\{\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}+\lambda_{2}({\mathbf{I}}-{\mathbf{Q}}_{2}{\mathbf{Q}}_{2}^{T})\right)_{-}\}.$
Similarly, for
@@
we need to ensure $\sigma_{1}\geq\max\{\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}\right)_{-}\}$ and $\sigma_{2}\geq\max\{\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}\right)_{-}\}$. ∎
Corollary 4.4 (One-sided updates for orthogonal ${\mathbf{A}}$).
If ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ and sketching matrix ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ has orthogonal rows, $\lambda\in[0,1]$ and the nonnegativity correction term $\sigma\geq\max\{\left({\mathbf{A}}^{T}{\mathbf{A}}\right)_{-}\}$, then the objective
$\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}+\sigma\|\mathbf{1}^{T}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}$ (4.5)
@@
${\mathbf{U}}$ $\leftarrow{\mathbf{U}}\circ\frac{{\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{X}}{\mathbf{V}}+\sigma\mathbf{1}\mathbf{1}^{T}{\mathbf{X}}{\mathbf{V}}}{(1-\lambda){\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}+\sigma\mathbf{1}\mathbf{1}^{T}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}+\lambda{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}}$ (4.6)
@@
Here, $\mathbf{1}=(1,\ldots,1)\in{\mathbb{R}}^{m}$. Note that Corollary 3.13 claims that the optimal solution for (4.5) results in a good solution for the uncompressed NMF problem as long as the original NMF error $\min\limits_{{\mathbf{U}},{\mathbf{V}}\geq 0}\|{\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}\|^{2}_{F}$ and $\|P^{\perp}_{\mathbf{A}}{\mathbf{X}}\|^{2}_{F}$ term are small.
@@
In the language of Theorem 4.1, “${\mathbf{X}}_{i}$” matrices are $\{{\mathbf{X}},\mathbf{0},{\mathbf{X}}\}$ and “${\mathbf{A}}_{i}$” matrices are $\{{\mathbf{A}},\sqrt{\lambda}({\mathbf{I}}-{\mathbf{A}}^{T}{\mathbf{A}}),\sqrt{\sigma}\mathbf{1}^{T}\}$ for $i=1,2,3$ respectively. One can see that
@@
These matrices are entry-wise nonnegative if $\sigma\geq\max\{\left({\mathbf{A}}^{T}{\mathbf{A}}\right)_{-}\}$ and $\lambda\in[0,1]$. Theorem 4.1 applies to justify that the objective (4.5) is nonincreasing under the updates (4.6). ∎
@@
Corollary 4.5 (One-sided updates for nonorthogonal sketching matrices).
@@ 4.3. Solving compressed problems with projected gradient descent
where $\alpha$ is the step size. For the sake of concreteness, we will give an example of the updates for one of our formulated objective functions. The projected gradient descent updates for the objective $\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}$ (as in Theorem 3.4) are
${\mathbf{U}}$ $\leftarrow\left({\mathbf{U}}-\alpha\left((1-\lambda){\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}+{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}-{\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{X}}{\mathbf{V}}\right)\right)_{+}$ (4.7)
@@
We can similarly derive updates for our other objective functions. A disadvantage of this method is that it possesses no guarantee of convergence or even a nonincreasing property. Empirically, we see (Figure 3 below) that on some datasets projected GD shows competitive performance and that its convergence is indeed not monotonic, unlike the convergence of the regularized compressed MU algorithm.
@@ 5. Experiments
5.1. Synthetic data: exact recovery from compressed measurements is achievable
Synthetic data has exactly low nonnegative rank $k=20$. Yet, the theoretical results suggest the possibility of exact recovery from the compressed data in two cases: (a) if two-sided compression was employed (Theorem 3.1), or (b) if the compression matrix is such that the projection of the data onto its orthogonal complement $P_{{\mathbf{A}}}^{\perp}{\mathbf{X}}$ is zero, for example, if the compression matrix is learned via the randomized rangefinder algorithm (one-sided data-adapted compression, Corollary 3.6). Two-sided compression requires twice more memory for the same sketch size but works with generic (data-oblivious) measurements that can be advantageous for various reasons as they do not require access to data or another pass over the data.
@@ 5.1. Synthetic data: exact recovery from compressed measurements is achievable
Figure 1. NMF recovery of the synthetic data with MU from full data (Uncompressed); from compressed data with one-sided data-adapted sketches using only $4\%$ of original memory (One-sided), and with two-sided Gaussian sketches using $8\%$ of original memory (Two-sided). All three methods achieve recovery within $10^{-3}$ relative error.
@@
In Figure 1, we compare nonnegative factors found by the MU from full data ${\mathbf{X}}$, from the one-sided data-adapted sketches, and from the two-sided oblivious (i.i.d. Gaussian) linear measurements. We take the target rank $r=20$ and the sketch size $k=20$ (so, the sketching matrices have the shape $20\times 1000$). In this case, the matrix ${\mathbf{X}}$ has one million elements whereas the total number of elements in ${\mathbf{A}}$ and ${\mathbf{X}}{\mathbf{A}}$ combined is only forty thousand. This represents a memory ratio of $4\%$ for the one-sided method. For the two sided method, the total number of elements in ${\mathbf{A}}_{1}$,${\mathbf{A}}_{2}$,${\mathbf{A}}_{1}{\mathbf{X}}$, and ${\mathbf{X}}{\mathbf{A}}_{2}$ is eighty thousand representing a memory ratio of $8\%$.
@@
We employ (compressed) MU algorithm as in Corollary 4.4 with $\lambda=0.1$ and Corollary 4.3 with $\lambda=0$. The parameter $\sigma$ is chosen minimal so that we have ${\mathbf{A}}{\mathbf{A}}^{T}\geq-\sigma$, and similarly for the ${\mathbf{A}}_{1}$, ${\mathbf{A}}_{2}$, $\sigma_{1}$ and $\sigma_{2}$ in the two-sided case. We see that the algorithm achieves a near exact fit, eventually reducing relative error below $10^{-3}$ which was our stopping criterion in this case. The one-sided method and the oblivious two-sided methods seem to be converging at a similar rate as the uncompressed method, albeit after a “burn-in” phase.
5.2. Synthetic data: effect of compression size
[…]
We show that on an “easy” synthetic problem oblivious one-sided measurements can be used: the compressed MU algorithm results in a good, although not perfect, recovery (Figure 2 (a,b)). The amount of limiting loss depending on the amount of compression. Then, the one-sided data-adapted compression and two-sided random compression attain exact reconstruction at different rates depending on the amount of compression (Figure 2 (c,d)). Not surprisingly, in every case, less compression leads to a faster or better convergence performance.

### v2

@@ 4.1. General convergence for sketched multiplicative updates
4.2. MU for solving regularized compressed problems
Now, we demonstrate how the general framework (20) applies to the compressed problems from Section 3.
Corollary 4 (Two-sided updates).
Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be a nonnegative matrix and ${\mathbf{A}}_{1}\in{\mathbb{R}}^{k\times m}$, ${\mathbf{A}}_{2}\in{\mathbb{R}}^{n\times k}$ are generic compression matrices such that ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}$ is invertible. Consider ${\mathbf{Q}}_{1}\in{\mathbb{R}}^{m\times r}$ and ${\mathbf{Q}}_{2}\in{\mathbb{R}}^{n\times r}$, the matrices which columns form orthonormal bases of the column space of ${\mathbf{X}}{\mathbf{A}}_{2}$ and the row space of ${\mathbf{A}}_{1}{\mathbf{X}}$ respectively. For any $\lambda_{1},\lambda_{2}\geq 0$, if
$\sigma_{1}$ $\geq\max\big(\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}\right)_{-},\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}+\lambda_{1}({\mathbf{I}}-{\mathbf{Q}}_{1}{\mathbf{Q}}_{1}^{T})\right)_{-}\big),$
$\sigma_{2}$ $\geq\max\big(\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}\right)_{-},\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}+\lambda_{2}({\mathbf{I}}-{\mathbf{Q}}_{2}{\mathbf{Q}}_{2}^{T})\right)_{-}\big),$
where $\max({\mathbf{M}}_{1},{\mathbf{M}}_{2})$ denotes maximum of the entries in all of the matrices ${\mathbf{M}}_{i}$, and $({\mathbf{M}})_{-}$ denotes negative part of the matrix, then the objective
@@ 4.2. MU for solving regularized compressed problems
$+\lambda_{1}\|P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}+\lambda_{2}\|{\mathbf{U}}{\mathbf{V}}^{T}P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}\|^{2}_{F}$
@@
Note that Theorem 1 and Corollary 2 claim that the optimal solution for (26) is the optimal solution for the uncompressed NMF problem if $\mathbf{X}$} has exactly nonnegative decomposition of the rank at most $k$.
@@
Consider the setting of Theorem 5, with the matrices
@@
$\{{\mathbf{B}}_{j}\}_{j=1,2,3}=\{{\mathbf{A}}_{2},\sqrt{\sigma_{2}}\mathbf{1}_{n},\sqrt{\lambda_{1}}P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}\}=\{{\mathbf{A}}_{2},\sqrt{\sigma_{2}}\mathbf{1}_{n},\sqrt{\lambda_{1}}({\mathbf{I}}-{\mathbf{Q}}_{1}{\mathbf{Q}}_{1}^{T})\},$
@@
Clearly, we have that the matrices ${\mathbf{X}}^{({\mathbf{A}})}_{i}$ and ${\mathbf{X}}^{({\mathbf{B}})}_{i}$ are nonnegative. Then to apply Theorem 5, we must check that $\sum_{i}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}$, $\sum_{i}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}{\mathbf{X}}^{({\mathbf{A}})}_{i}$, $\sum_{j}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T}$ and $\sum_{j}{\mathbf{X}}^{({\mathbf{B}})}_{j}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T}$ are entry-wise nonnegative. First, we calculate
@@
(24) $\sigma_{1}\geq\max\big(\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}+\lambda_{1}({\mathbf{I}}-{\mathbf{Q}}_{1}{\mathbf{Q}}_{1}^{T})\right)_{-}\big),$
$\sigma_{2}\geq\max\big(\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}+\lambda_{2}({\mathbf{I}}-{\mathbf{Q}}_{2}{\mathbf{Q}}_{2}^{T})\right)_{-}\big),$
that is, $\sigma_{1}$ and $\sigma_{2}$ upper-bound maximum elements in the respective matrices. Similarly, for
@@
we need to ensure
(25) $\sigma_{1}\geq\max\big(\left({\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}\right)_{-}\big)\;\text{ and }\;\sigma_{2}\geq\max\big(\left({\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}\right)_{-}\big).$
With $\sigma_{1}$ and $\sigma_{2}$ large enough to satisfy both (24) and (25), we conclude the proof of Corollary 4 as a special case of Theorem 5. ∎
Corollary 5 (One-sided updates for orthogonal $\mathbf{A}$}).
If ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ and sketching matrix ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ has orthogonal rows, $\lambda\in[0,1]$ and the nonnegativity correction term $\sigma\geq\max\big(\left({\mathbf{A}}^{T}{\mathbf{A}}\right)_{-}\big)$, then the objective
(26) $\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}+\sigma\|\mathbf{1}^{T}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}$
@@
(27) ${\mathbf{U}}$ $\leftarrow{\mathbf{U}}\circ\frac{{\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{X}}{\mathbf{V}}+\sigma\mathbf{1}\mathbf{1}^{T}{\mathbf{X}}{\mathbf{V}}}{(1-\lambda){\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}+\sigma\mathbf{1}\mathbf{1}^{T}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}+\lambda{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}}$
@@
Here, $\mathbf{1}=(1,\ldots,1)\in{\mathbb{R}}^{m}$. Note that Corollary 3 claims that the optimal solution for (26) results in a good solution for the uncompressed NMF problem as long as the original NMF error $\min\limits_{{\mathbf{U}},{\mathbf{V}}\geq 0}\|{\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}\|^{2}_{F}$ and $\|P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}$ term are small.
@@
In the language of Theorem 5, ${\mathbf{X}}_{i}$” matrices are $\{{\mathbf{X}},\mathbf{0},{\mathbf{X}}\}$ and ${\mathbf{A}}_{i}$” matrices are $\{{\mathbf{A}},\sqrt{\lambda}({\mathbf{I}}-{\mathbf{A}}^{T}{\mathbf{A}}),\sqrt{\sigma}\mathbf{1}^{T}\}$ for $i=1,2,3$ respectively. One can see that
@@
These matrices are entry-wise nonnegative if $\sigma\geq\max\big(\left({\mathbf{A}}^{T}{\mathbf{A}}\right)_{-}\big)$ and $\lambda\in[0,1]$. Theorem 5 applies to justify that the objective (26) is nonincreasing under the updates (27). ∎
@@
Corollary 6 (One-sided updates for nonorthogonal sketching matrices).
@@
To conclude, we note that all the considered sketched MU updates result in the computational cost of order $(n+m)kr$, in the natural case when $k,r\ll n,m$. At the same time, full MU update takes order nmrnmr operations per iteration. In particular, in the case when $n\sim m$, the quadratic per-iteration cost with respect to the size of the data becomes linear. The iterations of the sketched algorithms are indeed much faster, but they still could (and usually do) require more iterations until empirical convergence to a set threshold. In Appendix A we demonstrate this tradeoff empirically.
@@ 4.3. Solving compressed problems with projected gradient descent
where $\alpha$ is the step size. For the sake of concreteness, we will give an example of the updates for one of our formulated objective functions. The projected gradient descent updates for the objective $\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}$ (as in Theorem 2) are
(28) ${\mathbf{U}}$ $\leftarrow\left({\mathbf{U}}-\alpha\left((1-\lambda){\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}+{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}-{\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{X}}{\mathbf{V}}\right)\right)_{+}$
@@
We can similarly derive updates for our other objective functions. A disadvantage of GD with a constant step size is that it possesses no guarantee of convergence or even a nonincreasing property. We note that by choosing step sizes stepsize small enough, or adaptively, one can derive a gradient descent method with a nonincreasing guarantee, but for the sake of simplicity we do not consider this here. Empirically, we see (Figure 4 below) that on some datasets projected GD shows competitive performance and that its convergence is indeed not monotonic, unlike the convergence of the regularized compressed MU algorithm. We note that smaller step sizes in GD lead to slower convergence of the method, and the right trade-off can be dataset-dependent.
@@ 5. Experiments
All algorithms which we test require initial values for the factor matrices $\mathbf{U}$} and $\mathbf{V}$}. We note that the choice of initialization can significantly influence the convergence and overall behavior of NMF algorithms. To ensure fairness across experiments, we always initialize the matrices $\mathbf{U}$} and $\mathbf{V}$} randomly with all entries being independent standard log-normal random variables. We also fix the same random seed across comparisons.
@@
5.1. Exact recovery from compressed measurements is achievable
@@
Figure 1. NMF recovery of the synthetic data with MU from full data (Uncompressed); from compressed data with one-sided data-adapted sketches using only $4\%$ of original memory (One-sided), and with two-sided sketches using $8\%$ of original memory (Two-sided). In both cases, i.i.d. Gaussian sketching matrices are used.
All three methods achieve recovery within $10^{-3}$ relative error.
@@
To showcase the exact recovery, we focus on synthetic dataset that has an exact nonnegative factorization with $r=20$. In Figure 1, we compare nonnegative factors found by the MU from full data $\mathbf{X}$}, from the one-sided data-adapted sketches, and from the two-sided oblivious (i.i.d. Gaussian) linear measurements. We take the target rank $r=20$ and the sketch size $k=20$ (so, the sketching matrices have the shape $20\times 1000$). In this case, the matrix $\mathbf{X}$} has one million elements whereas the total number of elements in $\mathbf{A}$} and ${\mathbf{X}}{\mathbf{A}}$ combined is only forty thousand. This represents a memory ratio of $4\%$ for the one-sided method. For the two sided method, the total number of elements in ${\mathbf{A}}_{1}$${\mathbf{A}}_{2}$${\mathbf{A}}_{1}{\mathbf{X}}$, and ${\mathbf{X}}{\mathbf{A}}_{2}$ is eighty thousand representing a memory ratio of $8\%$.
@@
We employ (compressed) MU algorithm as in Corollary 5 with $\lambda=0.1$ and Corollary 4 with $\lambda=0$. The parameter $\sigma$ is chosen minimal so that we have ${\mathbf{A}}{\mathbf{A}}^{T}\geq-\sigma$, and similarly for the ${\mathbf{A}}_{1}$, ${\mathbf{A}}_{2}$, $\sigma_{1}$ and $\sigma_{2}$ in the two-sided case. We see that the algorithm achieves a near exact fit, eventually reducing relative error below $10^{-3}$ which was our stopping criterion in this case. The one-sided method and the oblivious two-sided methods seem to be converging at a similar rate as the uncompressed method, albeit after a “burn-in” phase.
5.2. Effect of compression size $k$ on recovery
In Figure 2, we compare the compressed multiplicative updates for different amounts of compression on the same synthetic dataset: the target rank $r=20$ and the sketch size $k$ varies.